OpenStudy (anonymous):

What is the solution set of x^2 + 5x - 5 = 0?

OpenStudy (anonymous):

\[x= [-5 \pm \sqrt{(5^2) - 4(1)(-5)} ] \div 2(1)\]

OpenStudy (anonymous):

so you will get x=\[[-5 + 3\sqrt{5} ] /2 \] and \[[−5-3\sqrt{5}] /2\]

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

btw sweet pic

OpenStudy (anonymous):

do you know the formula to find solutions which is \[[-b \pm \sqrt{b^2 - 4ac} ] /2a\] use this formula whenever u cant factorize the equation. a is the coefficient of x^2, b is the coefficient for x and c is the constant. lol thanks. :)

OpenStudy (anonymous):

ok thanks